OpenStudy (anonymous):

A baseball player hits a baseball into the outfield. The following equation models the ball's path. If not caught, when will the baseball land? Round to the nearest tenth second. h = -16t2 + 50t + 4

OpenStudy (goten77):

quadractict equasion well have \[(-50\pm \sqrt{50^{2}- (4\times -16 \times 4)}) / 2 \times -16\]

OpenStudy (goten77):

which becomes \[(-50 \pm \sqrt{2500 - (-256}) / -32\]